OpenStudy (jennychan12):

find lim x->2 (x^2-4)/(x+2) using limit definition

OpenStudy (anonymous):

lim(x->2){ x^2 - 4 / x + 2 } = lim(x->2){ (x - 2) (x + 2) / (x + 2) } = lim(x->2){ (x - 2) } = 2-2 = 0

OpenStudy (kainui):

Looks like you can just plug it in. The limit is equal to the function at that place since you're not dividing by zero.

OpenStudy (jennychan12):

i did it that way but my teacher want me to use the definition \[\frac{ \frac{ (x+h)^2 }{ s+h+2 } -\frac{ x^2-4 }{ x+2 }}{ h }\] i just need help simplifying...

OpenStudy (anonymous):

if it was like this : lim(x->2){ x^2 - 4 / x - 2 } = lim(x->2){ (x - 2) (x + 2) / (x - 2) } = lim(x->2){ (x + 2) } = 2+2 = 4

OpenStudy (kainui):

That's not the limit definition, that's the definition of a derivative.

OpenStudy (anonymous):

derivative or W H A T

OpenStudy (anonymous):